# A hybrid numerical method for elastic wave propagation in discontinuous media with complex geometry

### v1

@@ 3.1 The FD-DG interface and projection operators
To couple the FD and DG domains introduced in Section 2 we use the notion of norm compatible operators developed by Kozdon and Wilcox in [16]. The assumption of norm compatibility will prove to be crucial when deriving an energy estimate. In [16], the model problem was the wave equation in first order form. When applied to a second-order formulation with non-conforming FD-FD interfaces, however, this resulted in suboptimal convergence [22]. To recover optimal convergence, we therefore use the notion of projection pairs introduced in [3].
As in the previous section, let $\Gamma=\partial\Omega_{u}\cap\partial\Omega_{v}$ denote the interface. Moreover, let $\mathbf{x}_{u}$ be the uniform FD nodes along $\Gamma$, ordered left to right, and $\mathscr{T}_{\tilde{u}}$ the mesh whose element boundaries are defined by the nodes in $\mathbf{x}_{u}$. Similarly, we let $\mathscr{T}_{v}$ denote the edges in the DG mesh along the interface, and $\mathscr{T}_{g}$ the mesh with both FD and DG nodes as element boundaries. The setup of these grids are illustrated in Figure 1. For completeness, we also introduce the corresponding function spaces $\mathscr{F}_{\tilde{u}}$, $\mathscr{F}_{g}$ and $\mathscr{F}_{v}$ of piece-wise continuous polynomials of order $q$ over $\mathscr{T}_{\tilde{u}}$,$\mathscr{T}_{g}$ and $\mathscr{T}_{v}$ respectively.
FD $\mathbf{x}_{u}$$\mathscr{T}_{\tilde{u}}$dG $\mathscr{T}_{v}$lue $\mathscr{T}_{g}$
@@
Now let $f$ be a smooth function defined on an interval $J\subset\mathbb{R}$ containing $\Gamma$, with $\boldsymbol{f}_{u}$ representing its grid evaluation on $\mathbf{x}_{u}$. We introduce the projection operator $\Pi_{u2v}$, which maps the entries in $\boldsymbol{f}_{u}$ to the coefficients of a corresponding piece-wise continuous polynomial $\mathscr{P}f$ on $\mathscr{T}_{v}$. Let $\{\varphi^{\gamma}_{j}\}_{j=0}^{q}$ denote a set of local nodal basis functions in $\mathscr{F}_{v}$ associated with element $\gamma$ in $\mathscr{T}_{v}$, i.e. $\varphi_{j}^{\gamma}(x_{k})=\delta_{jk}$, where $\delta_{jk}$ denotes the Dirac-delta. Moreover, let $(\Pi_{u2v}\boldsymbol{f}_{u})^{\gamma}_{j}$ denote the projected value at node $j$ in element $\gamma$. Then, since the basis is nodal, $\mathscr{P}f(x)$ can be expanded in the element $\gamma$ as $\mathscr{P}f(x)=\sum_{j=0}^{q}(\Pi_{u2v}\boldsymbol{f})^{\gamma}_{j}\varphi_{j}^{\gamma}(x)$. Similarly, we can construct an operator $\Pi_{v2u}$, transforming the coefficients of the piece-wise polynomial back to the FD grid. As demonstrated in Figure 1, we do not force the nodes in $\mathbf{x}_{u}$ and $\mathscr{T}_{v}$ to align. To remedy this, we introduce the aforementioned glue grid, $\mathscr{T}_{g}$ having the nodes in $\mathbf{x}_{u}$ and $\mathscr{T}_{v}$ as element boundaries. To this end, we can first transform point-wise values from $\mathbf{x}_{u}$ to piece-wise continuous polynomials on $\mathscr{T}_{\tilde{u}}$ via the operator $\Pi_{u2\tilde{u}}$. Then we perform a basis change from $\mathscr{F}_{\tilde{u}}$ to $\mathscr{F}_{g}$ and construct a projection operator $\Pi_{\tilde{u}2g}$, followed by another basis change from $\mathscr{F}_{g}$ to $\mathscr{F}_{v}$ resulting in the operator $\Pi_{g2v}$. The product of these then forms our compound projection operators $\Pi_{u2v}=\Pi_{g2v}\Pi_{\tilde{u}2g}\Pi_{u2\tilde{u}}$ and $\Pi_{v2u}=\Pi_{\tilde{u}2u}\Pi_{g2\tilde{u}}\Pi_{v2g}$. In order to prove stability of the coupled scheme, we utilize the following notion of norm-compatible operators.
@@
We construct $\Pi_{u2\tilde{u}}$ such that the pair $\Pi_{u2\tilde{u}},\Pi_{\tilde{u}2u}$ is norm compatible. Then, as the remaining projection operators correspond to basis changes, they are norm compatible by construction. Since $\mathscr{F}_{\tilde{u}}$ and $\mathscr{F}_{v}$ are subspaces of the glue-space $\mathscr{F}_{g}$, following result is implied.
@@
as projection errors arising when projecting between $\mathbf{x}_{u}$ and $\mathscr{T}_{\tilde{u}}$. If we use a polynomial basis, it is natural to require that the errors vanish for polynomials up to a given order $q$. The projection operator will use a centered stencil in the interior, but a one-sided stencil near the boundary. Therefore, we must also distinguish between the boundary order of accuracy $q_{\partial}(\Pi_{u2v})$ and the interior order $q_{i}(\Pi_{u2v})$. Notably, using a DG method with basis of order $12p-$ 1 requires us to use $q_{i}(\Pi_{u2v})=2p-1$. It was however argued in [3] that norm compatibility requires the boundary order to be bounded as $q_{\partial}(\Pi_{u2v})\leq q_{i}(\Pi_{u2v})-1$. For a deeper discussion on the matter, see [3, 35].
The remaining projection operators $\Pi_{\tilde{u}2g},\Pi_{g2v},\Pi_{v2g}$ and $\Pi_{g2\tilde{u}}$ all correspond to basis transformations between function spaces. Since the basis functions all are of the same degree, the projection orders do not affect the order of accuracy. Thereby, the order of the overall projection operators $\Pi_{u2v}$ and $\Pi_{v2u}$ is completely determined by the accuracy of $\Pi_{u2\tilde{u}}$ and $\Pi_{\tilde{u}2u}$.
We shall follow the projection method outlined in [35] of projection pairs of good and bad operators $(\Pi^{g}_{u2v},\Pi^{b}_{v2u})$, $(\Pi^{b}_{u2v},\Pi^{g}_{v2u})$, where superscript $g$ denotes the good operator and $b$ the bad one, which also are norm compatible. The notion of good and bad relates to the orders of accuracy, which are $q_{\partial}(\Pi^{g})=q_{i}(\Pi^{g})-1=2p-2$ and $q_{\partial}(\Pi^{b})=q_{i}(\Pi^{b})-2=2p-3$ when using an FD method of order $2p$ and DG method with basis of order $12p-$ 1. Since the pairs are norm compatible, it is sufficient to set up a system of equations to solve for one operator from each pair and then use norm-compatibility to construct the remaining one. This procedure yields a set of free parameters, which in [35] were chosen by the authors to optimize accuracy. In this work, we instead optimize over these parameters with respect to both accuracy and spectrum, demonstrating a significant improvement as compared to choosing some arbitrary parameters.
@@ 3.2 Stability of the coupled method
We now turn to the DG semi-discretization and take inspiration of the form (8). However, we only consider the contribution to the fluxes of the edges in the set $\mathscr{E}_{h}^{\Gamma}$, consisting of edges having non-zero intersection with $\Gamma$. This is possible since the internal fluxes combine with the bulk terms analogously as in the pure DG case, and hence can be neglected. Moreover, for notational convenience, we rewrite the bulk in terms of the strain tensor $\varepsilon(\boldsymbol{v})=\frac{1}{2}(\nabla\boldsymbol{v}+\nabla\boldsymbol{v}^{T})$ and let $(\cdot,\cdot)_{\Omega_{v}}$ denote the standard $L^{2}(\Omega_{v})$ inner product. We then get the coupled semi-discretization as: find $\boldsymbol{v}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$ such that for each time $t\in(0,T]$ it holds
@@
$+\frac{1}{2}\sum_{\gamma\in\mathscr{E}_{h}^{\Gamma}}(\boldsymbol{v}_{\Gamma}-\boldsymbol{u}^{\gamma}_{\Gamma},\tau^{\gamma}(\boldsymbol{\varphi}))_{\gamma}+\frac{1}{2}\sum_{\gamma\in\mathscr{E}_{h}^{\Gamma}}(\tau^{\gamma}(\boldsymbol{v})-\tau^{\gamma}(\boldsymbol{u}),\boldsymbol{\varphi})_{\gamma}$
@@
for every $\boldsymbol{\varphi}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$. Here we have introduced the notation $\boldsymbol{v}_{\Gamma}^{\flat}=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{b}\mathcal{P}_{v2u}^{g}\boldsymbol{v}_{\Gamma})^{\gamma}_{i}\boldsymbol{\varphi}_{i}^{\gamma}$, and $\boldsymbol{u}_{\Gamma}^{\gamma}=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{g}\boldsymbol{u}_{\Gamma})^{\gamma}_{i}\boldsymbol{\varphi}_{i}^{\gamma}$ for the component of $\boldsymbol{u}_{\Gamma}$ projected from $\mathbf{x}_{u}$ onto $\gamma\in\mathscr{E}_{h}^{\Gamma}$, where $\{\boldsymbol{\varphi}^{\gamma}_{i}(x)\}_{i=0}^{q}$ again denotes the local nodal basis on the edge $\gamma$. Analogously we introduce the projected quantities $\tau^{\gamma}(\boldsymbol{u})=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{b}\mathcal{T}_{u}\boldsymbol{u})_{i}^{\gamma}\boldsymbol{\varphi}^{\gamma}_{i}$ and $\tilde{\boldsymbol{u}}^{\gamma}_{\Gamma}=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{b}\boldsymbol{u}_{\Gamma})^{\gamma}_{i}\boldsymbol{\varphi}_{i}^{\gamma}$, where $\tau^{\gamma}(\boldsymbol{u})$ corresponds to the FD traction projected onto the edge $\gamma$ and $\tilde{\boldsymbol{u}}_{\Gamma}$ is a projection onto $\gamma$ by the bad projection operator. We recall that the DOFS are ordered in a way consistent with the order of $\mathbf{x}_{u}$. With this in place. we can set to prove the main theoretical contribution of the paper.
@@
Let $\chi_{1}=2\mu_{v}+\lambda_{v}$ and $\chi_{2}=2\mu_{u}+\lambda_{u}$, assume that the operators $\mathcal{D}_{y}$, $\mathcal{D}_{yy}$ are fully compatible and the projection operators between the FD and DG mesh are norm compatible, then the semi-discretizations (6) and (12) form an energy conserving scheme under the choices $\alpha_{u}=\alpha_{v}\geq\frac{1}{4C}$ and $\beta_{u}=\beta_{v}\geq\frac{1}{4C^{\prime}}$ for some constants $C,C^{\prime}>0$ depending on the quality of $\mathscr{K}_{h}$.
@@
We begin by considering the discretization (12). In order to write the semi-dicretization on a form compatible with (LABEL:eq:FDdiscretization) we need to introduce the full-size one-dimensional mass matrix $M_{\Gamma}$ along the interface and $\mathcal{M}_{\Gamma}=I_{2}\otimes M_{\Gamma}$ such that $(\boldsymbol{\varphi}_{i},\boldsymbol{\varphi}_{j})_{\Gamma}=\boldsymbol{\varphi}_{i}^{T}\mathcal{M}_{\Gamma}\boldsymbol{\varphi}_{j}$ for $\boldsymbol{\varphi}_{i},\boldsymbol{\varphi}_{j}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$. We also need an operator $\mathcal{T}_{v}$ for the DG discretization such that for a given edge $\gamma\in\mathscr{E}_{h}$ we can represent the traction as $\tau^{\gamma}(\boldsymbol{v})=\sum_{i=0}^{q}(\mathcal{T}_{v}\boldsymbol{v})^{\gamma}_{i}\boldsymbol{\varphi}_{i}$. Then, letting $\boldsymbol{\varphi}=\boldsymbol{v}_{t}$ we can write the semi-discretization as
@@
This allows us to rewrite the bad projection operators in terms of the good ones, which upon gathering terms only proportional to $\boldsymbol{v}$ by themselves and terms containing both $\boldsymbol{u},\boldsymbol{v}$ by themselves yields
@@
(($\star$)( ⋆ ) $=-\boldsymbol{u}_{t}^{T}\mathcal{H}_{x}\mathcal{T}_{u}\boldsymbol{u}-\boldsymbol{u}_{t}^{T}\big{(}\mathcal{L}_{2}\mathcal{D}_{x}^{T}\mathcal{H}\mathcal{D}_{y}+\mathcal{L}_{2}^{T}\mathcal{D}_{y}^{T}\mathcal{H}\mathcal{D}_{x}+\mathcal{L}_{1}\mathcal{D}_{x}^{T}\mathcal{H}_{y}\mathcal{D}_{x}$ (17)
$+\mathcal{L}_{3}\mathcal{D}_{y}^{T}\mathcal{H}_{x}\mathcal{D}_{y}\big{)}\boldsymbol{u}=-\boldsymbol{u}_{t}^{T}\mathcal{H}_{x}\mathcal{T}_{u}\boldsymbol{u}-\boldsymbol{u}_{t}^{T}\mathcal{U}\boldsymbol{u},$
where we for notational convenience introduce the matrix $\mathcal{U}$ satisfying $\mathcal{U}=\mathcal{U}^{T}$ Combining this with the last two lines in (15) we can once more gather $\boldsymbol{u}$ and $\boldsymbol{v}$ terms as well as writing everything in terms of good projection operators by using (13), resulting in
@@
To prove stability it now remains to show that the combination of (18) and (14) results in a change of energy which is non-positive. Upon comparing the two equations it is immediate that by setting $\alpha_{u}=\alpha_{v}$ and $\beta_{v}=\beta_{u}$ we can add the two equations to get
$\frac{1}{2}\frac{d}{dt}\big{(}\rho_{u}\boldsymbol{u}_{t}^{T}\mathcal{H}\boldsymbol{u}_{t}+(\rho_{v}\boldsymbol{v}_{t},\boldsymbol{v}_{t})_{\Omega_{v}}\big{)}=-\frac{1}{2}\frac{d}{dt}\big{(}U_{1}+U_{2}\big{)},$ (19)
where we have introduced the quantities
@@
We set to show that $U_{1}\geq 0$. By completing the square we can rewrite it as
@@
Note that the second row is non-negative under the choice of $\chi_{1}$ and $\alpha_{v}$ positive, hence it is sufficient to consider the first row. This will follow from Lemma 9 and Lemma 2.2 since $2\mu_{v}\|\varepsilon(\boldsymbol{v})\|^{2}_{\Omega_{v}}+\lambda_{v}\|\nabla\cdot\boldsymbol{v}\|^{2}_{\Omega_{v}}=\int_{\Omega_{v}}\sigma(\boldsymbol{v}):\varepsilon(\boldsymbol{v})dA$, yielding
@@
where it again is sufficient to bound the terms on the top row. Analogous to the previous calculation, we can utilize an inverse estimate to bound the term $\boldsymbol{u}^{T}\mathcal{U}\boldsymbol{u}$. In this case, the estimate is given by
$\boldsymbol{u}^{T}\mathcal{H}\boldsymbol{u}\geq C^{\prime}h_{u}\boldsymbol{u}^{T}\mathcal{H}_{x}\boldsymbol{u},$
where $C^{\prime}>0$ depends on the smallest eigenvale of the matrix $\mathcal{H}_{y}$. This allows us, after some algebra, to write the first row of (23) in the following form

### v2

@@ 3.1 The FD-DG interface and projection operators
To couple the FD and DG domains introduced in Section 2 we use the notion of norm compatible operators developed by Kozdon and Wilcox in [18]. The assumption of norm compatibility will prove to be crucial when deriving an energy estimate. In [18], the model problem was the wave equation in first order form. When applied to a second-order formulation with non-conforming FD-FD interfaces, however, using the operators from [18] resulted in suboptimal convergence [24]. To recover optimal convergence, we therefore use the notion of projection pairs introduced in [3].
As in the previous section, let $\Gamma=\partial\Omega_{u}\cap\partial\Omega_{v}$ denote the interface. Moreover, let $\mathbf{x}_{u}$ be the uniform FD nodes along $\Gamma$, ordered left to right, and $\mathscr{T}_{\tilde{u}}$ the mesh whose element boundaries are defined by the nodes in $\mathbf{x}_{u}$. Similarly, we let $\mathscr{T}_{v}$ denote the edges in the DG mesh along the interface, and $\mathscr{T}_{g}$ the mesh with both FD and DG nodes as element boundaries. The setup of these grids are illustrated in Figure 1. For completeness, we also introduce the corresponding function spaces $\mathscr{F}_{\tilde{u}}$, $\mathscr{F}_{g}$ and $\mathscr{F}_{v}$ of piecewise continuous polynomials of order $q$ over $\mathscr{T}_{\tilde{u}}$,$\mathscr{T}_{g}$ and $\mathscr{T}_{v}$ respectively.
@@
Now let $f$ be a smooth function defined on $\Gamma$, with $\boldsymbol{f}_{u}$ representing its grid evaluation on $\mathbf{x}_{u}$. We introduce the projection operator $\Pi_{u2v}$ such that $\Pi_{u2v}\boldsymbol{f}_{u}$ contains the coefficients of a piecewise continuous polynomial $\mathscr{P}f$ on $\mathscr{T}_{v}$ that approximates $f$. Let $\{\varphi^{\gamma}_{j}\}_{j=0}^{q}$ denote a set of local nodal basis functions in $\mathscr{F}_{v}$ associated with element $\gamma$ in $\mathscr{T}_{v}$, i.e. for a node $x_{k}$ in the element $\gamma$, $\varphi_{j}^{\gamma}(x_{k})=\delta_{jk}$, where $\delta_{jk}$ denotes the Dirac-delta. Moreover, let $(\Pi_{u2v}\boldsymbol{f}_{u})^{\gamma}_{j}$ denote the projected value at node $j$ in element $\gamma$. Then, since the basis is nodal, $\mathscr{P}f(x)$ can be expanded in the element $\gamma$ as $\mathscr{P}f(x)|_{\gamma}=\sum_{j=0}^{q}(\Pi_{u2v}\boldsymbol{f})^{\gamma}_{j}\varphi_{j}^{\gamma}(x)$. Similarly, we can construct an operator $\Pi_{v2u}$, transforming the coefficients of the piecewise polynomial back to the FD grid. As demonstrated in Figure 1, we do not force the nodes in $\mathbf{x}_{u}$ and $\mathscr{T}_{v}$ to align. To remedy this, we introduce the aforementioned glue grid, $\mathscr{T}_{g}$ having the nodes in $\mathbf{x}_{u}$ and $\mathscr{T}_{v}$ as element boundaries. To this end, we can first transform pointwise values from $\mathbf{x}_{u}$ to piecewise continuous polynomials on $\mathscr{T}_{\tilde{u}}$ via the operator $\Pi_{u2\tilde{u}}$. The operator $\Pi_{u2\tilde{u}}$ is therefore in general not exact. However, note that both $\mathscr{F}_{v}$ and $\mathscr{F}_{\tilde{u}}$ are subspaces of $\mathscr{F}_{g}$. We can therefore construct an exact projection operator $\Pi_{g2\tilde{u}}$ from $\mathscr{F}_{g}$ to $\mathscr{F}_{\tilde{u}}$, followed by another similar operator $\Pi_{g2v}$ from $\mathscr{F}_{g}$ to $\mathscr{F}_{v}$. The product of these then forms our compound projection operators $\Pi_{u2v}=\Pi_{g2v}\Pi_{\tilde{u}2g}\Pi_{u2\tilde{u}}$ and $\Pi_{v2u}=\Pi_{\tilde{u}2u}\Pi_{g2\tilde{u}}\Pi_{v2g}$. In order to prove stability of the coupled scheme, we utilize the following notion of norm-compatible operators.
@@
We construct $\Pi_{u2\tilde{u}}$ such that the pair $\Pi_{u2\tilde{u}},\Pi_{\tilde{u}2u}$ is norm compatible. Then, as the remaining projection operators correspond to projections between subspaces, they are norm compatible by construction. Since $\mathscr{F}_{\tilde{u}}$ and $\mathscr{F}_{v}$ are subspaces of the glue-space $\mathscr{F}_{g}$, the following result is implied.
@@
as projection errors arising when projecting between $\mathbf{x}_{u}$ and $\mathscr{T}_{\tilde{u}}$. If we use a polynomial basis, it is natural to require that the errors vanish for polynomials up to a given order $q$. The projection operator will use a centered stencil in the interior, but a one-sided stencil near the boundary. Therefore, we must also distinguish between the boundary order of accuracy $q_{\partial}(\Pi_{u2v})$ and the interior order $q_{i}(\Pi_{u2v})$. Notably, using a method of interior order $q_{i}(\Pi_{u2v})$ makes the projection errors vanish for all polynomials up to order $q_{i}(\Pi_{u2v})-1$. Moreover, it was shown in [24] that for norm compatibility to hold between two FD grids, the overall orders of accuracy on the boundary of the projection operators when using a quadrature rule of $2p$ should be related by $q_{\partial}(\Pi_{u2v})+q_{\partial}(\Pi_{v2u})\leq 2p+1$. For a deeper discussion, see [3, 35].
The remaining projection operators $\Pi_{\tilde{u}2g},\Pi_{g2v},\Pi_{v2g}$ and $\Pi_{g2\tilde{u}}$ all correspond to transformations between function spaces which can be seen as subspaces of $\mathscr{F}_{g}$. Since the basis functions all are of the same degree, the projection orders do not affect the order of accuracy. Thereby, the order of the overall projection operators $\Pi_{u2v}$ and $\Pi_{v2u}$ is completely determined by the accuracy of $\Pi_{u2\tilde{u}}$ and $\Pi_{\tilde{u}2u}$.
We shall follow the projection method outlined in [35] of projection pairs of good and bad operators $(\Pi^{g}_{u2\tilde{u}},\Pi^{b}_{\tilde{u}2u})$, $(\Pi^{b}_{u2\tilde{u}},\Pi^{g}_{\tilde{u}2u})$, where superscript $g$ denotes the good operator and $b$ the bad one, which also are norm compatible. The notion of good and bad relates to the orders of accuracy. From the discussion above on the bound of the order of accuracy we let the orders for the good and bad orders be given by $q_{\partial}(\Pi^{g})=p+1,q_{i}(\Pi^{g})=2p$ and $q_{\partial}(\Pi^{b})=p,q_{i}(\Pi^{b})=2p$ when using an FD method of order $2p$ and DG method with basis of order $p+1$. Since the pairs are norm compatible, it is sufficient to set up a system of equations to solve for one operator from each pair and then use norm-compatibility to construct the remaining one. This procedure yields a set of free parameters, which in [35] were chosen by the authors to optimize accuracy. In this work, we instead optimize over these parameters with respect to both accuracy and spectrum, demonstrating a significant improvement as compared to choosing some arbitrary parameters.
@@ 3.2 Stability of the coupled method
We now turn to the DG semi-discretization and take inspiration of the form (8). However, we only consider the contribution to the fluxes of the edges in the set $\mathscr{E}_{h}^{\Gamma}$, consisting of edges having non-zero intersection with $\Gamma$. This is possible since the internal fluxes combine with the bulk terms analogously as in the DG case, and hence can be neglected. Moreover, for notational convenience, we rewrite the bulk in terms of the strain tensor $\varepsilon(\boldsymbol{v})=\frac{1}{2}(\nabla\boldsymbol{v}+\nabla\boldsymbol{v}^{T})$ and let $(\cdot,\cdot)_{\Omega_{v}}$ denote the standard $L^{2}(\Omega_{v})$ inner product. We then get the coupled semi-discretization as: find $\boldsymbol{v}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$ such that for each time $t\in(0,T]$ it holds
@@
$+\frac{1}{2}\sum_{\gamma\in\mathscr{E}_{h}^{\Gamma}}(\boldsymbol{v}_{\Gamma}-\boldsymbol{u}^{\gamma}_{\Gamma},\boldsymbol{\tau}^{\gamma}(\boldsymbol{\varphi}))_{\gamma}+\frac{1}{2}\sum_{\gamma\in\mathscr{E}_{h}^{\Gamma}}(\boldsymbol{\tau}^{\gamma}(\boldsymbol{v})-\boldsymbol{\tau}^{\gamma}(\boldsymbol{u}),\boldsymbol{\varphi})_{\gamma}$
@@
for every $\boldsymbol{\varphi}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$. Here we have introduced the notation $\boldsymbol{v}_{\Gamma}^{\flat}=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{b}\mathcal{P}_{v2u}^{g}\boldsymbol{v}_{\Gamma})^{\gamma}_{i}\boldsymbol{\varphi}_{i}^{\gamma}$, and $\boldsymbol{u}_{\Gamma}^{\gamma}=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{g}\boldsymbol{u}_{\Gamma})^{\gamma}_{i}\boldsymbol{\varphi}_{i}^{\gamma}$ for the component of $\boldsymbol{u}_{\Gamma}$ projected from $\mathbf{x}_{u}$ onto $\gamma\in\mathscr{E}_{h}^{\Gamma}$, where $\{\boldsymbol{\varphi}^{\gamma}_{i}(x)\}_{i=0}^{q}$ again denotes the local nodal basis on the edge $\gamma$. Analogously we introduce the projected quantities $\boldsymbol{\tau}^{\gamma}(\boldsymbol{u})=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{b}\mathcal{T}_{u}\boldsymbol{u})_{i}^{\gamma}\boldsymbol{\varphi}^{\gamma}_{i}$ and $\tilde{\boldsymbol{u}}^{\gamma}_{\Gamma}=\sum_{i=0}^{q}(\mathcal{P}_{u2v}^{b}\boldsymbol{u}_{\Gamma})^{\gamma}_{i}\boldsymbol{\varphi}_{i}^{\gamma}$, where $\boldsymbol{\tau}^{\gamma}(\boldsymbol{u})$ corresponds to the FD traction projected onto the edge $\gamma$ and $\tilde{\boldsymbol{u}}_{\Gamma}$ is a projection onto $\gamma$ by the bad projection operator. We recall that the nodes are ordered in a way consistent with the order of $\mathbf{x}_{u}$. With this in place. we can set to prove the main theoretical contribution of the paper.
@@
Let $\chi_{1}=2\mu_{v}+\lambda_{v}$ and $\chi_{2}=2\mu_{u}+\lambda_{u}$, where the Lamé parameters are positive. Assume that the operators $\mathcal{D}_{y}$, $\mathcal{D}_{yy}$ are fully compatible and the projection operators between the FD and DG mesh are norm compatible, then the semi-discretizations (6) and (12) form an energy conserving scheme under the choices $\alpha_{u}=\alpha_{v}\geq\frac{1}{4C}$ and $\beta_{u}=\beta_{v}\geq\frac{1}{4C^{\prime}}$ for some constants $C,C^{\prime}>0$, with $C$ depending on the quality of $\mathscr{K}_{h}$ and $C^{\prime}$ on the order of the SBP operators used.
@@
We begin by considering the discretization (12). In order to write the semi-dicretization on a form compatible with (LABEL:eq:FDdiscretization) we need to introduce the full-size one-dimensional mass matrix $M_{\Gamma}$ along the interface and $\mathcal{M}_{\Gamma}=I_{2}\otimes M_{\Gamma}$ such that $(\boldsymbol{\varphi}_{i},\boldsymbol{\varphi}_{j})_{\Gamma}=\boldsymbol{\varphi}_{i}^{T}\mathcal{M}_{\Gamma}\boldsymbol{\varphi}_{j}$ for $\boldsymbol{\varphi}_{i},\boldsymbol{\varphi}_{j}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$. We also need an operator $\mathcal{T}_{v}$ for the DG discretization such that for a given edge $\gamma\in\mathscr{E}_{h}$ we can represent the traction as $\boldsymbol{\tau}^{\gamma}(\boldsymbol{v})=\sum_{i=0}^{q}(\mathcal{T}_{v}\boldsymbol{v})^{\gamma}_{i}\boldsymbol{\varphi}_{i}$. Then, letting $\boldsymbol{\varphi}=\boldsymbol{v}_{t}$ we can write the semi-discretization as
@@
This allows us to rewrite the bad projection operators in terms of the good ones. Gathering terms only containing $\boldsymbol{v}$ by themselves and terms with both $\boldsymbol{u},\boldsymbol{v}$ by themselves yields
@@
(($\star$)( ⋆ ) $=-\boldsymbol{u}_{t}^{T}\mathcal{H}_{x}\mathcal{T}_{u}\boldsymbol{u}-\boldsymbol{u}_{t}^{T}\big{(}\mathcal{L}_{2}\mathcal{D}_{x}^{T}\mathcal{H}\mathcal{D}_{y}+\mathcal{L}_{2}^{T}\mathcal{D}_{y}^{T}\mathcal{H}\mathcal{D}_{x}+\mathcal{L}_{1}\mathcal{D}_{x}^{T}\mathcal{H}\mathcal{D}_{x}$ (17)
$+\mathcal{L}_{3}\mathcal{D}_{y}^{T}\mathcal{H}\mathcal{D}_{y}\big{)}\boldsymbol{u}=-\boldsymbol{u}_{t}^{T}\mathcal{H}_{x}\mathcal{T}_{u}\boldsymbol{u}-\boldsymbol{u}_{t}^{T}\mathcal{U}\boldsymbol{u},$
where we for notational convenience introduce the matrix
$\mathcal{U}=\mathcal{L}_{2}\mathcal{D}_{x}^{T}\mathcal{H}\mathcal{D}_{y}+\mathcal{L}_{2}^{T}\mathcal{D}_{y}^{T}\mathcal{H}\mathcal{D}_{x}+\mathcal{L}_{1}\mathcal{D}_{x}^{T}\mathcal{H}\mathcal{D}_{x}+\mathcal{L}_{3}\mathcal{D}_{y}^{T}\mathcal{H}\mathcal{D}_{y}$
, which by the virtue of $\mathcal{L}_{1},\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ being block diagonal satisfies $\mathcal{U}=\mathcal{U}^{T}$ Combining this with the last two lines in (15) we can once more gather $\boldsymbol{u}$ and $\boldsymbol{v}$ terms as well as writing everything in terms of good projection operators by using (13), resulting in
@@
To prove stability it now remains to show that the combination of (18) and (14) results in a non-positive quantity. Upon comparing the two equations it is immediate that by setting $\alpha_{u}=\alpha_{v}$ and $\beta_{v}=\beta_{u}$ we can add the two equations to get
$\frac{1}{2}\frac{dE}{dt}=-\frac{1}{2}\frac{d}{dt}\big{(}U_{1}+U_{2}\big{)},$ (19)
where we have introduced the quantities $E=\rho_{u}\boldsymbol{u}_{t}^{T}\mathcal{H}\boldsymbol{u}_{t}+(\rho_{v}\boldsymbol{v}_{t},\boldsymbol{v}_{t})_{\Omega_{v}}$,
@@
If we then can show that $U_{1},U_{2}\geq 0$ it follows that $\tilde{E}=E+U_{1}+U_{2}$ corresponds to a semi-discrete energy which is conserved. We set to show that $U_{1}\geq 0$. By completing the square we can rewrite it as
@@
Note that the second row is non-negative under the choice of $\chi_{1}$ and $\alpha_{v}$ positive, hence it is sufficient to consider the first row. This will follow from Lemma 2.1 and Lemma 2.2 since $2\mu_{v}\|\varepsilon(\boldsymbol{v})\|^{2}_{\Omega_{v}}+\lambda_{v}\|\nabla\cdot\boldsymbol{v}\|^{2}_{\Omega_{v}}=\int_{\Omega_{v}}\sigma(\boldsymbol{v}):\varepsilon(\boldsymbol{v})dA$, yielding
@@
where it again is sufficient to bound the terms in the first row of $U_{2}$. A key step to obtain a bound is the positivity property from section 4.2 in [2] for the SBP norm, which is an analogue to the trace inequality Lemma 2.2 used on the DG side. In our case, the estimate is given by
$\boldsymbol{u}^{T}\mathcal{H}\boldsymbol{u}\geq C^{\prime}h_{u}\boldsymbol{u}^{T}\mathcal{H}_{x}\boldsymbol{u},$ (24)
where $C^{\prime}>0$ depends on the value of the quadrature matrix $H$ at the boundary and is given for SBP operators of order $62p=2,4,$ 6 in Table 1.
Order $(2p)$ $C^{\prime}$
2 0.5
4 0.3542
6 0.3159
Table 1: Value of the constant $C^{\prime}$ used in the estimate (24) for SBP operators of order 2,4 and 6.
[…]